OpenStudy (anonymous):

Will fan and medal!!!! Please help!!! Prove that the following relationship is true: nCr + nCr + 1 = n + 1Cr + 1 Use the least common denominator method.

OpenStudy (anonymous):

\[\large\begin{align*}{}_nC_r+{}_nC_{r+1}&=\frac{n!}{r!(n-r)!}+\frac{n!}{(r+1)!(n-r-1)!}\\\\ &=\frac{n!\color{red}{(r+1)}}{\color{red}{(r+1)}r!(n-r)!}+\frac{n!\color{red}{(n-r)}}{(r+1)!\color{red}{(n-r)}(n-r-1)!}\\\\ &=\frac{n!\color{red}{(r+1)}}{(r+1)!(n-r)!}+\frac{n!\color{red}{(n-r)}}{(r+1)!(n-r)!}\\\\ &=\frac{n!\left(r+1+n-r\right)}{(r+1)!(n-r)!}\\\\ &=\frac{n!\left(n+1\right)}{(r+1)!(n-r)!}\\\\ &=\frac{\left(n+1\right)!}{(r+1)!((n+1)-(r+1))!} \end{align*}\]

OpenStudy (anonymous):

@sithsandgiggles THANK YOU SOOOOO MUCHHHH